OpenStudy (jmartinez638):

Sketch the graph of a function f(x) for which f(x) > 0, f’(x) > 0, and f’’(x) > 0 on the entire domain of f ?

OpenStudy (jmartinez638):

@SolomonZelman

OpenStudy (solomonzelman):

Hello!

OpenStudy (jmartinez638):

Heya! :D

OpenStudy (solomonzelman):

I check polynomial wouldn't work. It would have t be even to make f(x)>0, but f'(x) would be >0 at least for some x.

OpenStudy (solomonzelman):

it is probably a square root

OpenStudy (zarkon):

a very famous function works

OpenStudy (solomonzelman):

famous function?

OpenStudy (solomonzelman):

I get something like \(\color{#0000ff }{ \displaystyle f(x)=\sqrt{x^3+3} }\)

OpenStudy (zarkon):

a function all should know

OpenStudy (solomonzelman):

yes pliz

OpenStudy (zarkon):

\[\Large f(x)=e^x\]

OpenStudy (solomonzelman):

Oh...

OpenStudy (solomonzelman):

yes of course :))

OpenStudy (solomonzelman):

I thought maybe \(\color{#0000ff }{ \displaystyle f(x)=\sqrt{x+3}+3 }\) fits

OpenStudy (solomonzelman):

assuming that f'(x) doesn't have to be defined everywhere where f(x) is defined

OpenStudy (zarkon):

the second derivative of that function is not positive

OpenStudy (solomonzelman):

yes, the power....

OpenStudy (solomonzelman):

we can make it positive... if we make a power bigger but of same nature, perhaps

OpenStudy (zarkon):

with the cubed...then yes

OpenStudy (solomonzelman):

\(\color{#0000ff }{ \displaystyle f(x)=\sqrt{(x+3)^5} +4 }\)

OpenStudy (solomonzelman):

yes or cubed... i did 5th power just to secure myself

OpenStudy (jmartinez638):

:| A lot is happening xD

OpenStudy (solomonzelman):

do you have any questions about anthing in (or outside) the thread?

OpenStudy (jmartinez638):

Not anything in this thread, anymore, but I have a few questions about other things in calc entirely xD

OpenStudy (solomonzelman):

ok:)

OpenStudy (solomonzelman):

I hope I am denoting things correctly.

OpenStudy (solomonzelman):

I think I have extracted the general solution for this... \(\color{#0000ff }{ \displaystyle y=a^{\Large x} ;\quad a>1}\) \(\color{#0000ff }{ \displaystyle y=\sqrt[\Large 2b]{(x+5)^{\Large 3b+2m}};\quad b,m\in\mathbb{N} }\)

OpenStudy (solomonzelman):

don't want to get to complicated tho... you might as well take e^x ...

OpenStudy (solomonzelman):

good luck

OpenStudy (jmartinez638):

Ok xD thanks guys

OpenStudy (solomonzelman):

yw